f(x)= 3(x-1)(x+2) / (x-4) (x+2)

f) state horizontal asymptotes

g) sketch of graph


1
Expert's answer
2021-09-16T00:46:49-0400
"f(x)=\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}"

"x-4\not=0, x+2\not=0"


"Domain: (-\infin, -2)\cap (-2, 4)\cap (4, \infin)"

(f)


"\lim\limits_{x\to-\infin}f(x)=\lim\limits_{x\to-\infin}\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}"

"=\lim\limits_{x\to-\infin}\dfrac{3(x-1)}{x-4}=\dfrac{3}{1}=3"


"\lim\limits_{x\to\infin}f(x)=\lim\limits_{x\to\infin}\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}"

"=\lim\limits_{x\to\infin}\dfrac{3(x-1)}{x-4}=\dfrac{3}{1}=3"

Horizontal asymptote: "y=3"


(g)


"\lim\limits_{x\to-2^-}f(x)=\lim\limits_{x\to-2^-}\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}"

"=\lim\limits_{x\to-2^-}\dfrac{3(x-1)}{x-4}=\dfrac{3(-2-1)}{-2-4}=\dfrac{3}{2}"



"\lim\limits_{x\to-2^+}f(x)=\lim\limits_{x\to-2^+}\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}"


"=\lim\limits_{x\to-2^+}\dfrac{3(x-1)}{x-4}=\dfrac{3(-2-1)}{-2-4}=\dfrac{3}{2}"


"\lim\limits_{x\to-2^-}f(x)=\dfrac{1}{2}=\lim\limits_{x\to-2^+}f(x)=>\lim\limits_{x\to-2}f(x)=\dfrac{1}{2}"

The function "f(x)" has a removable discontinuity at "x=-2."



"\lim\limits_{x\to4^-}f(x)=\lim\limits_{x\to4^-}\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}"

"=\lim\limits_{x\to4^-}\dfrac{3(x-1)}{x-4}=-\infin"





"\lim\limits_{x\to4^+}f(x)=\lim\limits_{x\to4^+}\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}""=\lim\limits_{x\to4^+}\dfrac{3(x-1)}{x-4}=\infin"

Vertical asymptote: "x=4"


"x=0, f(0)=\dfrac{3(0-1)(0+2)}{(0-4)(0+2)}=\dfrac{3}{4}"

Point "(0, \dfrac{3}{4})" is "y" -intercept.


"y=0, 0=\dfrac{3(x-1)(x+2)}{(x-4)(x+2)}=>x=1"

Point "(1, 0)" is "x" -intercept.